# Two-Player Zero-Sum Games with Bandit Feedback

### v1

@@ 4 ETC-TPZSG Algorithm and Regret Analysis
2:$S_{A}$: set of action pairs in the game matrix $A$triangleright $S_{A}$ includes action pairs $(i,j)$
@@
7:$\sigma^{2}$: subgaussian variance factor
8:Initialize: $\hat{A}=[0]^{m\times l}$triangleright $\hat{A}$ is estimated game matrix
9:In round $t=0,1,2,...,mlk$; $\triangleright$ Exploration Phase
@@
11:        Update $\hat{A}$ using 3
12:In round $t=mlk+1,mlk+2,...,T$; $\triangleright$ Committing Phase
13:        Play an equilibrium $(i^{*},j^{*})$ satisfying the following property:
$\hat{A}(i,j^{*})\leq\hat{A}(i^{*},j^{*})\leq\hat{A}(i^{*},j)$
@@
We provide the proof in Appendix B. The regret comprises two main components, the exploration phase and the loss due to misidentifying the NE. Thus, it is crucial to choose an appropriate exploration time $k$ that balances sufficient exploration and efficient exploitation. Choosing $k$ too large leads to unnecessary exploration while setting it too small increases the risk of making suboptimal decisions. Hence, careful tuning of $k$ is essential to minimize overall regret.
Let assume there are two action pairs $(i_{1},j)$ and $(i_{2},j)$ in the game such that row player has two actions and column player has one action. In addition, let suppose the NE is at $(i_{1},j)$. Then, we have $\Delta=A(i_{1},j)-A(i_{2},j)$ and we can write the regret simply as
$R_{T}^{*}\leq k\Delta+(T-2k)\Delta\exp{\Big{(}-\frac{k\Delta^{2}}{16\sigma^{2}}\Big{)}}\leq k\Delta+T\Delta\exp{\Big{(}-\frac{k\Delta^{2}}{16\sigma^{2}}\Big{)}}.$ (15)
Taking the first derivative with respect to $k$ and solve it, we get the following for the exploration time:
@@
where the first term $T\Delta$ is worst-case regret and for the other, we combine the regret we obtain in (15) with the value of exploration time in (16). Then, focusing on $\max\Big{\{}0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big{\}}$, we obtain that $\Delta=\frac{4\sigma}{\sqrt{T}}$ is a critical point. If we put it in the regret bound (17), we have the following:
$R_{T}^{*}\leq\Delta+c\sqrt{T}$
where $c>0$ is a some constant. If $\Delta\leq 1$, it becomes
$R_{T}^{*}\leq 1+c\sqrt{T}.$ (18)
Bounds like (18) are referred to as instance-independent since it only depends on the time horizon $T$ not on the game instance. In the zero-sum game setting, our analysis shows that we get a regret bound comparable to the standard bandit setting in Lattimore_Szepesvari_2020 , indicating that the ETC-based algorithm performs effectively in adversarial environments and aligns with known results in the literature.
5 ETC-TPZSG-AE Algorithm and Regret Analysis
If a pure NE exists at $(i^{*},j^{*})$, then it must satisfy the condition given in (2). For the algorithm in the section, we make use of the concept of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), which approximately satisfies the standard NE condition. Specifically, an $\varepsilon$-NE satisfies the NE conditions within a tolerance level $\varepsilon$, making it a suitable criterion in learning settings where exact equilibrium identification is hard.
Based on this criterion, we introduce an elimination strategy in Algorithm 2, inspired by auer2010ucb , to reduce unnecessary exploration of clearly suboptimal action pairs. Specifically, action pairs that do not satisfy the $\varepsilon$-NE property are eliminated from further play, which enables a more efficient learning process by concentrating exploration on approximately optimal action pairs.
Definition 5.1.
@@ 5 ETC-TPZSG-AE Algorithm and Regret Analysis
For the exploration time $k$ of the algorithm, let just go back to the regret bound in (14) and assume that there exists some $\Delta^{*}$ and $\Delta$ such that $\Delta^{*}\geq\Delta_{ij}^{*}$ and $\Delta\leq\Delta_{ij}$ for all $(i,j)\in S_{A}$. Then, when we put them into the regret bound and take its derivative with respect to $k$, we obtain a similar result as in (16); thus, as the exploration time, clearly we can use the following:
$k=\Big{\lceil}\frac{16\sigma^{2}}{\Delta^{2}}\ln{\Big{(}\frac{\Delta^{2}T}{16\sigma^{2}}\Big{)}}\Big{\rceil}$ (20)
@@
Since the true suboptimality gaps $\Delta$ are unknown, the algorithm adopts a decreasing $\Delta$ approach across rounds. It is motivated by the intuition that the remaining action pairs become closer to the true NE as suboptimal action pairs are progressively eliminated. Consequently, the suboptimality gaps among the remaining action pairs are expected to decrease over time. In other words, as the rounds progress, we expect that the players get closer to true NE, which means they tend to repeat better action pairs more often. We can afford to use smaller $\Delta$ values to explore action pairs near the NE more thoroughly, since smaller $\Delta$ leads to a longer exploration time $k$.
@@
The algorithm schedules $\Delta_{ij}$ across rounds using $\hat{\Delta}_{t}$, which guides both exploration time and the tolerance level for the $\varepsilon$-NE condition. When $\hat{\Delta}_{t}$ is large, action pairs are played fewer times and more action pairs are kept during elimination. As $\hat{\Delta}_{t}$ decreases, the algorithm focuses exploration on pairs closer to equilibrium, having already eliminated clearly suboptimal ones.
@@
In order to perform updates over rounds, an initial estimate is required. When the payoffs are bounded within a known interval, such as $[0,1]$, it is common practice to initialize the estimated suboptimality gap $\hat{\Delta}_{t}$ with value 1, representing the maximum possible difference between arm rewards as in auer2010ucb . However, in our setting, the payoffs are assumed to be $\sigma$-subgaussian, which implies that there is no strict upper bound on the suboptimality gaps, we only have $\Delta_{ij}\geq 0$. Since no empirical estimates are available at the initialization step, it is not feasible to use a bound for the expected value of the maximum of $\sigma$-subgaussian random variables to guide the choice of $\hat{\Delta}_{t}$. To address this, we initialize $\hat{\Delta}_{t}$ with $4\sigma$, which provides a practical starting point for the analysis.
@@
Moreover, we use $\hat{\Delta}_{t}=2^{-t+2}\sigma$ in each round $t$ which takes values from 00 to $\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$. Regarding $\Delta\approx\sqrt{(16\sigma^{2}/k)\ln{(\Delta^{2}T/16\sigma^{2})}}$ from (20) so $\ln{(\Delta^{2}T/16\sigma^{2})}\geq 0$, which gives us $\Delta\geq\sqrt{16\sigma^{2}/T}$. Then, using $2^{-t+2}\sigma\geq\sqrt{16\sigma^{2}/T}$ and after some calculations, we obtain $t\leq\frac{1}{2}\log_{2}T$. The division by $e$ is a technical adjustment, which makes the bound safer in the analysis; otherwise the number of play would be zero in the last round.
@@
If we assume that a unique pure NE exists in the game as before, a well estimated game matrix $\hat{A}$ should also exhibit the properties necessary to support the equilibrium. Specifically, $\hat{A}$ must satisfy the conditions required for the $\varepsilon$-NE in (19) to hold. Based on this principle, the Algorithm 2 systematically eliminates action pairs that fail to satisfy this condition, as such pairs cannot be near to the equilibrium. By reducing the set of action pairs to include only the pairs that meet this criterion, it ensures that the remaining action pairs are the only ones that could potentially be near the NE; thus, it enables to focus on the most relevant action pairs. Consequently, we expect that the algorithm converges to an accurate approximation of the NE, even in the presence of estimation errors in $\hat{A}$.
@@
2:$S_{A}$: set of action pairs in the game matrix $A$triangleright $S_{A}$ includes action pairs $(i,j)$
@@
6:$\sigma^{2}$: subgaussian variance factor
7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$triangleright $\hat{A}$ is estimated game matrix
@@
10:            If $\mathbf{|S_{t}|>1}:$triangleright Exploration Phase
11:              Explore each action pair $(i,j)$ in $S_{t}$ $k_{t}=\Big{\lceil}\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}\Big{\rceil}$ times

### v2

@@ 4 ETC-TPZSG Algorithm
[…]
@@
[…]
8:Initialize: $\hat{A}=[0]^{m\times l}$ $\triangleright$ $\hat{A}$ is estimated game matrix
9:In round t=0,1,2,,mlkt=0,1,2,...,mlk; $\triangleright$ Exploration Phase
@@
11:    Update $\hat{A}$ using (4)
12:In round t=mlk+1,mlk+2,,Tt=mlk+1,mlk+2,...,T; $\triangleright$ Committing Phase
13:    Play an equilibrium $(i^{*},j^{*})$ satisfying the following
14:    property:
$\hat{A}(i,j^{*})\leq\hat{A}(i^{*},j^{*})\leq\hat{A}(i^{*},j),\quad\forall i\in S_{x},j\in S_{y}$
@@
Proof Sketch.
The total regret consists of two components: the regret from exploration and the regret from exploitation. During the exploration phase, each action pair is played $k$ times. In the exploitation phase, the algorithm plays according to the pure equilibria, where we account for the probability that the pure NE is not correctly identified. In other words, in the exploitation phase, we analyze the loss incurred from playing with a misidentified NE. Thus, the expected number of plays for any suboptimal action pair $(i,j)$ can be written by the following:
$\mathbb{E}[n_{ij,T}]=k+(T-Nk)\ \mathbb{P}((i,j)\text{ is identified as NE})$
Once this probability is bounded, the proof is complete. A detailed derivation is provided in Appendix B. ∎
As mentioned above, the regret comprises two main components, the exploration phase and the loss due to misidentifying the NE. Thus, it is crucial to choose an appropriate exploration time $k$ that balances sufficient exploration and efficient exploitation. Choosing $k$ too large leads to unnecessary exploration, while setting it too small increases the risk of making suboptimal decisions. Hence, careful tuning of $k$ is essential to minimize overall regret.
Let assume there are two action pairs $(i_{1},j)$ and $(i_{2},j)$ in the game such that the row player has two actions and the column player has one action. In addition, let suppose the NE is at $(i_{1},j)$. Then, we have $\Delta=A(i_{1},j)-A(i_{2},j)$ and we can write the regret simply as
$R_{T}^{*}$ $\leq k\Delta+(T-2k)\Delta\exp{\Big(-\frac{k\Delta^{2}}{16\sigma^{2}}\Big)}$ (15)
$\leq k\Delta+T\Delta\exp{\Big(-\frac{k\Delta^{2}}{16\sigma^{2}}\Big)}.$ (16)
Taking the first derivative with respect to $k$ and solving it, we get the following for the exploration time:
@@
where the first term $T\Delta$ is the worst-case regret and for the other, we combine the regret we obtain in (16) with the value of the exploration time in (17). Then, focusing on $\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}$, we find that $\Delta=\frac{4\sigma}{\sqrt{T}}$ is a critical point. If we put it in the regret bound (18), we have the following:
$R_{T}^{*}\leq\Delta+c\sqrt{T}$ (19)
where $c>0$ is some constant. If $\Delta\leq 1$, it becomes $R_{T}^{*}\leq 1+c\sqrt{T}$, which refers to an instance-independent bound since it depends only on the time horizon $T$ not on the game instance. In the zero-sum game setting, our analysis shows that we get a regret bound similar to the standard bandit setting in [20], indicating that the ETC-based algorithm performs effectively in adversarial environments.
5 ETC-TPZSG-AE Algorithm
In this section, we extend Algorithm 1 by incorporating adaptive elimination, called ETC-TPZSG-AE. We use the concept of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), which approximately satisfies the standard NE condition, to eliminate action pairs. Specifically, if a pure NE exists at $(i^{*},j^{*})$, then it must satisfy the condition given in (3) while a $\varepsilon$-NE satisfies the NE conditions within a tolerance level $\varepsilon$, making it a suitable criterion in learning settings where exact equilibrium identification is challenging.
Based on this criterion, we introduce an elimination method in Algorithm 2, inspired by [6], to reduce unnecessary exploration of clearly suboptimal action pairs. Specifically, action pairs that do not satisfy the $\varepsilon$-NE property are eliminated from further play, which enables a more efficient learning process by concentrating exploration on approximately optimal action pairs.
Definition 5.
@@ 5 ETC-TPZSG-AE Algorithm
For the exploration time of the algorithm, let just go back to the regret bound in (14) and assume that there exist some $\Delta^{*}$ and $\Delta$ such that $\Delta^{*}\geq\Delta_{ij}^{*}$ and $\Delta\leq\Delta_{ij}$ for all $(i,j)\in S_{A}$. Then, when we put them into the regret bound and take its derivative with respect to $k$, we obtain a result similar to that of (17).
@@
Since the true suboptimality gaps $\Delta$ are unknown, the algorithm adopts a decreasing $\Delta$ approach over rounds. It is motivated by the intuition that the remaining action pairs become closer to the true NE as suboptimal action pairs are progressively eliminated. Consequently, the suboptimality gaps among the remaining action pairs are expected to decrease over time. In other words, as the rounds progress, we expect that the players get closer to true NE, which means that they tend to repeat better action pairs more often. We can afford to use smaller $\Delta$ values to explore action pairs near the NE more thoroughly, since a smaller $\Delta$ leads to a longer exploration time $k$.
@@
The algorithm schedules $\Delta_{ij}$ across rounds using $\hat{\Delta}_{t}$, which guides both the exploration time and the tolerance level for the $\varepsilon$-NE condition. When $\hat{\Delta}_{t}$ is large, the action pairs are played fewer times and more action pairs are kept during elimination. As $\hat{\Delta}_{t}$ decreases, the algorithm focuses exploration on pairs closer to equilibrium and eliminates clearly suboptimal ones. Thus, as the exploration time, we can clearly use the following:
$k_{t}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}\Big\rceil$ (21)
@@
In order to perform updates over rounds, an initial estimate is required. When payoffs are bounded within a known interval, such as $[0,1]$, it is common practice to initialize the estimated suboptimality gap $\hat{\Delta}_{t}$ with value 1, representing the maximum possible difference between arm rewards as in [6]. However, in our setting, the payoffs are assumed to be $\sigma$-subgaussian, which implies that there is no strict upper bound on the suboptimality gaps, we only have $\Delta_{ij}\geq 0$. Since no empirical estimates are available at the initialization step, it is not feasible to use a bound for the expected value of the maximum of $\sigma$-subgaussian random variables to guide the choice of $\hat{\Delta}_{t}$. To address this, we initialize $\hat{\Delta}_{t}$ with $4\sigma$, which provides a practical starting point for the analysis.
@@
We set $\hat{\Delta}_{t}=2^{-t+2}\sigma$ in each round $t$, where $t$ ranges from 0 to $\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$. We have $\hat{\Delta}_{t}\approx\sqrt{(16\sigma^{2}/k_{t})\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}}$ from (21), which implies $\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}\geq 0$. It gives us $\hat{\Delta}_{t}\geq\sqrt{16\sigma^{2}/T}$ and our initialization of $\hat{\Delta}_{t}$ remains consistent with this observation. Then, using $2^{-t+2}\sigma\geq\sqrt{16\sigma^{2}/T}$ and after some calculations, we obtain $t\leq\frac{1}{2}\log_{2}T$. The division by $e$ is a technical adjustment, which makes the bound safer in the analysis; otherwise, the number of plays would be zero in the last round.
@@
Since we assume that a unique pure NE exists in the game, a well estimated game matrix $\hat{A}$ should also exhibit the properties necessary to support the equilibrium. Specifically, $\hat{A}$ must satisfy the conditions required for the $\varepsilon$-NE in (20) to hold. Based on this principle, Algorithm 2 systematically eliminates action pairs that fail to satisfy this condition, as such pairs cannot be near to the equilibrium. By reducing the set of action pairs to include only the pairs that meet this criterion, it ensures that the remaining action pairs are the only ones that could potentially be near the NE; thus, it enables to focus on the most relevant action pairs. Consequently, we expect that the algorithm converges to an accurate approximation of the NE, even in the presence of estimation errors in $\hat{A}$.
Theorem 2.
where $\lambda\geq 4\sigma\sqrt{e/T}$, and $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
Proof Sketch.
To analyze the regret of Algorithm 2, at first we define $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ as the earliest round satisfying $\hat{\Delta}_{t}<\Delta_{ij}/2$ and then we consider four distinct cases. First, we analyze the suboptimal action pairs which are not eliminated in round $t_{ij}$ or before. Next, we examine the suboptimal pairs that are eliminated in the round $t_{ij}$ or earlier, allowing us to bound the number of plays with suboptimal action pairs. The third case examines the scenario where the optimal action pair is eliminated in $t_{ij}$. Finally, we consider the case in which a suboptimal action pair remains as a unique one in the game (discussed in detail in Appendix C). ∎
@@
[…]
@@
[…]
7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
@@
10:     If $\mathbf{|S_{t}|>1}:$ $\triangleright$ Exploration Phase
11:      Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times, defined in (21)
12:      Update $\hat{A}$ using (4)
13:     Else: $\triangleright$ Committing Phase
14:      Play with a unique action pair in $S_{t}$ until step $T$
15:    Action Pair Elimination:
$\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
17:    Update:
18:     $S_{t+1}$ by "Action Pair Elimination" step
Algorithm 2 ETC-TPZSG-AE
Theorem 3.
$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}+\lambda T\Big).$ (23)
Proof.
Although we have a different notion of regret, this does not affect the underlying structure of the regret analysis. This ensures that the previous theoretical bounds can be extended using $\Delta_{ij}^{*}\leq\Delta_{ij}$, then the result follows. ∎
We simplify the regret bound to $O(\frac{\log(T\Delta^{2})}{\Delta})$, as the logarithmic term dominates since by setting $\lambda=\sqrt{16\sigma^{2}e/T}$, we ensure that the term $\lambda T$ is bounded by $\sqrt{16\sigma^{2}eT}$, which is at most $\frac{e}{\Delta_{ij}}$ when $\Delta_{ij}\leq\lambda$. Our results are consistent with the regret bound derived by [6].
Compared to our regret bound $O(\frac{\log(T\Delta^{2})}{\Delta})$ with the bound $O(\frac{\log_{+}(T\Delta^{2})}{\Delta})$ in [18], we observe that our study introduces a more refined notion of the suboptimality gap $\Delta$. Specifically, in [18], $\Delta$ refers to the gap between pure NE and the action chosen when the opponent plays the action in equilibrium. Moreover, our regret bound is tighter, as their formulation involves the square of the sum of suboptimality gaps, while ours uses the sum of squares of suboptimality gaps within the logarithmic factor. Moreover, our elimination algorithm specifically incorporates the extended pure NE property through $\varepsilon$, which also allows us to explain why we achieve a tighter regret bound for a TPZSG setting where both players have a pure strategy.
As a result, we have $R_{T}^{*}\leq R_{T}$, where $R_{T}^{*}$ evaluates its performance against the NE, capturing the cumulative loss relative to the equilibrium strategy, while the regret analysis of $R_{T}$ allows us to consider the losses of both players with respect to their best individual responses in each round, providing a broader and more flexible notion of regret. This outcome aligns with the interpretation that the expected regret against the NE establishes a more specific comparison, and thus, it naturally provides a tighter upper bound on the expected regret.
6 ETC-TPZSG-AE-NUE Algorithm
In this section, we extend the adaptive elimination algorithm by introducing a variant, ETC-TPZSG-AE-NUE. It uses a non-uniform exploration method The non-uniform exploration in these algorithms is based on players’ strategies, which indirectly reflect the probability of being NE. The key difference from the previous algorithm is that Algorithm 3 explores action pairs in proportion to the strategies of the players derived from the estimated payoff matrix. Thus, it is employed to reduce the number of plays for suboptimal action pairs more effectively.
As mentioned in Section 3, the strategies of the row and column players are defined as probability distributions $p$ and $q$, respectively, over their corresponding action sets. In round $t$, using the estimated game matrix, we denote the estimated strategy of the row player by $\hat{p}_{t}$ and the estimated strategy of the column player by $\hat{q}_{t}$. Since the true game matrix is assumed to have the pure NE, we expect these estimated strategies to converge to the pure strategies over time.
The estimated strategies of the players serve as a basis to determine the exploration time in our proposed algorithms. Furthermore, the probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}_{i,t}\hat{q}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$. Consequently, in Algorithm 3, the exploration time for each action pair $(i,j)$ is proportional to $\hat{P}_{ij,t}$, allowing action pairs with higher estimated probabilities to be played more frequently.
We incorporate the strategies of the players $\hat{p}_{t}$ and $\hat{q}_{t}$ together with $k_{t}$. Specifically, in each round $t$, this involves estimating the optimal strategies for each player by solving a linear system using the estimated payoff matrix $\hat{A}$ as follows:
$\begin{aligned} &\\ &\hat{P}_{t}=\hat{p}_{t}\hat{q}_{t}\\ &\hat{p}_{t}=\max_{p}\min_{q}p^{T}\hat{A}q\\ &\hat{q}_{t}=\min_{q}\max_{p}p^{T}\hat{A}q\\ \end{aligned}\qquad\begin{aligned} &\text{Subject to:}\\ &\sum_{i\in S_{x}}\hat{p}_{i,t}=1,\quad\hat{p}_{i,t}\geq 0,\ \forall i\in S_{x}\\ &\sum_{j\in S_{y}}\hat{q}_{j,t}=1,\quad\hat{q}_{j,t}\geq 0,\ \forall j\in S_{y}\\ &\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=\sum_{i,j}\hat{p}_{i,t}\hat{q}_{j,t}=1\end{aligned}$ (24)
Thus, the exploration time is given by the following:
$k_{ij,t}=k_{t}+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t}}{\hat{\Delta}_{t}^{2}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}\Big\rceil$ (25)
where $k_{t}$ is defined in (21). Thus, the algorithm explores each action pair at least $k_{t}$ times. Since players’ strategies are expected to converge to pure ones over time, the Algorithm 3 progressively focuses on playing the optimal action pair more. This ensures that action pairs with higher estimated probabilities are explored more frequently, while those with lower probabilities are selected less often. Thus, these algorithms achieve a more efficient balance between exploration and exploitation by reducing unnecessary plays on suboptimal action pairs, which accelerates convergence toward optimal action pair.
We note that we set $t=\left\lfloor\frac{1}{4}\log_{2}\frac{T}{e}\right\rfloor$ as the last round because unlike the Algorithm 2, which assigns an exploration time of $k_{t}$, the current algorithms allow exploration up to $2k_{t}$ times, ensuring that each action pair is explored at least $k_{t}$ times. As a result, due to this extended exploration phase, the number of rounds is reduced by half.
1:Input:
[…]
3:$m$: number of actions for row player
4:$l$: number of actions for column player
5:$T$: time horizon $\triangleright$ $1\leq ml\ll T$
[…]
7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$, $\hat{A}=[0]^{m\times l}$ and $P_{ij,0}=\frac{1}{ml},\forall i,j$
8:In round $t=0,1,2,...,\left\lfloor\frac{1}{4}\log_{2}\frac{T}{e}\right\rfloor$;
9:    Action Pair Selection:
10:     If $\mathbf{|S_{t}|>1}:$ $\triangleright$ Exploration Phase
11:      Explore each $(i,j)$ in $S_{t}$ $k_{ij,t}$ times, defined in (25)